OpenStudy (anonymous):

Prove (cos^2x-sin^2x+sin^4x)/(sin^4x) = cot^4x

OpenStudy (anonymous):

Trying to solve csc^4x-2csc^2x+1=cot^4x...this is as far as I got.

OpenStudy (anonymous):

is it...\[\frac{ \cos ^{2}x-\sin ^{2}x+\sin ^{4}x }{ \sin ^{4}x }=\cot ^{4}x\]

OpenStudy (anonymous):

Yes...unless I made a mistake along the way.

OpenStudy (anonymous):

we can start factoring sin^2x...

OpenStudy (anonymous):

all expression of (1-sin^2x) = cos^2x....

OpenStudy (anonymous):

oops i can't finish it... i have to go, but if you'll just apply identity 1 (my last entry)... you'll arrive at... cos^4x/sin^4x = cot^4x.... i'm sorry i have to go....

OpenStudy (anonymous):

Thanks I guess. I know I have to end up a (cosx/sinx)^4

OpenStudy (anonymous):

we have a moderator viewing he can finish it....

OpenStudy (anonymous):

afk?

OpenStudy (anonymous):

ok i'm back... i miss my service....

OpenStudy (anonymous):

sorry that you missed it then.

OpenStudy (anonymous):

\[\frac{ \cos ^{2}x-\sin ^{2}x(1-\sin ^{2}x) }{ \sin ^{4}x }=\frac{ \cos ^{2}x-\sin ^{2}x \cos ^{2}x }{ \sin ^{4}x }\]

OpenStudy (anonymous):

factor out cos^2x....

OpenStudy (anonymous):

\[\frac{ \cos ^{2}x(1-\sin ^{2}x) }{ \sin ^{4}x }=\frac{ \cos ^{2}x \cos ^{2}x }{ \sin ^{4}x }=\frac{ \cos ^{4}x }{ \sin ^{4}x }\]

OpenStudy (anonymous):

\[\frac{ \cos ^{4}x }{ \sin ^{4}x }=\cot ^{4}x\]

OpenStudy (anonymous):

that completes the proof.... :-)

OpenStudy (anonymous):

I've been on this for 2 hours :< Thank you very much.

OpenStudy (anonymous):

you're welcome... anytime... :)